# Are Statistical Methods Obsolete in the Era of Deep Learning?

### v1

@@ Abstract
In the era of AI, neural networks have become increasingly popular for modeling, inference, and prediction, largely due to their potential for universal approximation. With the proliferation of such deep learning models, a question arises: are leaner statistical methods still relevant? To shed insight on this question, we employ the mechanistic nonlinear ordinary differential equation (ODE) inverse problem as a testbed, using physics-informed neural network (PINN) as a representative of the deep learning paradigm and manifold-constrained Gaussian process inference (MAGI) as a representative of statistically principled methods. Through case studies involving the SEIR model from epidemiology and the Lorenz model from chaotic dynamics, we demonstrate that statistical methods are far from obsolete, especially when working with sparse and noisy observations. On tasks such as parameter inference and trajectory reconstruction, statistically principled methods consistently achieve lower bias and variance, while using far fewer parameters and requiring less hyperparameter tuning. Statistical methods can also decisively outperform deep learning models on out-of-sample future prediction, where the absence of relevant data often leads overparameterized models astray. Additionally, we find that statistically principled approaches are more robust to accumulation of numerical imprecision and can represent the underlying system more faithful to the true governing ODEs.
@@ 1 Introduction
The advancement of deep neural network models in the last fifteen years has profoundly altered the scientific landscape of estimation, prediction and decision making, from the early success of image recognition (Krizhevsky et al., 2012; He et al., 2016), to the success of self-learning of board games (Silver et al., 2017), to machine translation (Wu et al., 2016), to generative AI (Ho et al., 2020), and to the success of protein structure prediction (Jumper et al., 2021), among many other developments. In many of these successes, there are no well-established mechanistic models to describe the underlying problem (for example, we do not fully understand how human brains translate from one language to another). As such, it is conceivable that such successes are attributable to deep neural networks’ remarkable capabilities for universal function approximation. In contrast, the hand-crafted models that existed before deep neural networks (such as n-gram models (Katz, 1987; Brown et al., 1992; Bengio et al., 2000)) were too restricted to offer satisfactory approximation.
@@
How well do deep neural network models work when there are well-established mechanistic models (as in physical sciences, where decades of theoretical and experimental endeavor have yielded highly accurate mechanistic models in many cases) — in particular, how do the inference and prediction results of deep neural network models compare to more statistical approaches in the presence of reliable mechanistic models — is an interesting question. The current article focuses on this comparison.
@@
In physical and natural sciences, the mechanistic models are often described by ordinary differential equations (ODEs), which are often referred to as dynamical systems. Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo, 2006; Mueller and Siltanen, 2012), arises in many scientific studies, where advances in experimental and data collection techniques have allowed researchers to track dynamical systems closer to real-time (Xie, 2015). Although ODE models are parametric, their inference has long challenged researchers, as most nonlinear ODEs require numerical solutions, which can be expensive for many practical applications (Mueller and Siltanen, 2012).
@@
In this article, we compare the neural network approach to the statistical approach on their inference of dynamical systems. Both approaches were proposed to address the numerical challenge of (nonlinear) dynamical systems. Specifically, we compare the performance of the recently developed physics-informed neural network (PINN) (Raissi et al. (2017b, 2019); Lu et al. (2021)), which represents the neural network approach to the ODE inference problem, to that of manifold-constrained Gaussian process inference (MAGI) (Yang et al., 2021), which represents a principled Bayesian statistical approach to the ODE inference. In this comparison, we investigate three different aspects of statistical inference from noisy and sparse observations: (i) how well can the methods recover the model parameters (including quantities of scientific interest defined by the parameters), (ii) how well can the methods recover the system’s time-trajectories, and (iii) how well can the methods predict the future state of the system given the observations from a short time window. In the comparison, for a comprehensive understanding we also consider the case of missing system components, where some system components are entirely unobserved; such cases often arise in real experiments due to various (technical or experimental) limitations of data collection.
@@
Our study demonstrates that MAGI is generally more robust and reliable than PINN, particularly in sparse and noisy data environments that are common in laboratory sciences. In these environments, researchers often face fewer than 100 observation data points with noise levels ranging from 5%, 15% to higher. We found that for the SEIR model, when all system components are observed, the performance of PINN and MAGI is comparable. However, when missing components are present, MAGI clearly outperforms PINN. The contrast becomes even more pronounced with the chaotic Lorenz system, where PINN struggles to handle the complexity of the system, while MAGI continues to produce accurate inference results. Our results indicate that in the era of deep learning, statistically principled methods are still very relevant, especially when working with sparse and noisy data. Our study underscores the importance of proper probabilistic modeling in capturing the underlying dynamics for inference and prediction when one has limited data.
@@ 1.1 Statistical inference of dynamical systems
where the vector $\bm{x}(t)$ represents the system’s state/status at time $t$, and $\bm{\theta}$ is the model parameters. For example, in neuroscience, ODEs underpin models of neuron signal transmission (FitzHugh, 1961; Nagumo et al., 1962). In ecology, predator-prey population dynamics uses ODEs (Lotka, 1932) to capture interactions between species over time. In systems biology, ODEs map complex gene regulatory networks, enabling insight into cellular responses (Hirata et al., 2002). In epidemiology, ODE models are used to track disease progression and evaluate public health interventions (Shaman and Karspeck, 2012).
Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo, 2006; Mueller and Siltanen, 2012), includes estimating the model parameters, inferring/recovering the time-trajectories of the system, and making predictions of the underlying system at future times from noisy observations collected at various time points. The difficulty of inferring dynamical system models lies in the computational costs associated with numerically solving the ODEs, which has long challenged the applied mathematics community (Mueller and Siltanen, 2012).
Utilizing the approximating power of neural networks, the physics-informed neural network (PINN) was recently introduced (Raissi et al. (2017b, 2019)); it uses a neural network as a function approximator to the solution of the ODE, thus bypassing numerical integration, where the loss function of the neural network incorporates the ODE dynamics. Section 2 describes PINN in detail.
Parallel to the development of PINN, there are also efforts in the statistical community for the inference of dynamical systems. One competitive method is the manifold-constrained Gaussian process inference (MAGI) (Yang et al., 2021), which assigns a Gaussian process prior on the system components $\bm{x}(t)$ and explicitly conditions it to ensure that the derivatives of $\bm{x}(t)$ satisfy the ODE specification. This approach enables MAGI to also bypass the need for numerical integration. Section 2 also describe MAGI in detail.
@@ 1.2 Other approaches
When the function $f$ in the ODE (Equation (1)) is nonlinear, determining $\bm{x}(t)=\mathbb{X}_{\bm{x}_{0},\bm{\theta}}(t)$ given the initial conditions $\bm{x}_{0}$ and the parameter $\bm{\theta}$ often requires a numerical integration method, such as the fifth-order Runge–Kutta (Lapidus and Seinfeld, 1971) $\hat{\mathbb{X}}^{RK5}$. Given access to numerical integration, the inverse problem of estimating $\bm{\theta}$ from an experimental/observation dataset $\{y(\tau_{i}):i=1,2,\ldots,N\}$, where $y(\tau_{i})=X(\tau_{i})+\epsilon(\tau_{i})$ represents a noisy observation of the system at time $\tau_{i}$, can be cast as an optimization problem:
@@
Before the recent development of PINN and MAGI, various attempts have been made to mitigate this computational difficulty. In statistics, a cornerstone contribution that uses splines as a surrogate for $\bm{x}(t)$ was introduced in Ramsay et al. (2007). In constrained optimization, surrogate models are also proposed to bypass numerical solvers for ODE inference. Hard constraints for satisfying ODEs are imposed on collocation points. For example, Gaussian process surrogates are used as convenient interpolation tools with easy computation of the hard constraints (Raissi et al., 2017a; Spitieris and Steinsland, 2023; Oates et al., 2019; Chen et al., 2021). The hard constraints, however, are artificially imposed on the Gaussian process rather than a principled Bayesian derivation through conditioning (Cockayne et al., 2017; Graepel, 2003; Chen et al., 2022).
Other statistical methods (Müller and Timmer, 2002), especially the Bayesian model calibration (Kennedy and O’Hagan, 2001)), have also gained popularity (Chiachio-Ruano et al., 2021; Stuart, 2010; Stuart and Teckentrup, 2018). Many of these methods, however, still cannot completely bypass the numerical integration steps (Marzouk et al., 2007; Lan et al., 2016; Li and Marzouk, 2014; Santner et al., 2019). Those that bypass the numerical integration are mostly two-stage methods (Bär et al., 1999; Müller and Timmer, 2004; Rai and Tripathi, 2019), where the ODE solution and its derivatives are first reconstructed from observation data using basis function expansion, such as a polynomial basis (Franke and Schaback, 1998) or splines (Bär et al., 1999). Then, the differential equation parameters are estimated using regression from the first-step approximation results. Although simple to implement, two-stage methods are not reliable when the data is sparse and noisy (Müller and Timmer, 2004): they often require ad hoc follow-up approximations, such as Picard linearization or moment-matching Gaussian distribution for non-linear equations (Magnani et al., 2022; Xun et al., 2013; Zhou et al., 2022).
@@ 2.1 PINN: a neural network (NN) approach
Raissi et al. (2019) suggests using a neural network (“NN”) $N(t)$ as a function approximator of the system components $\bm{x}(t)$. Their approach incorporate the ODE physics information into the neural network loss function, popularizing the concept of a physics-informed neural network (PINN), with its loss function shown below:
@@
[…]
@@ 2.2 MAGI: a statistical Bayesian approach
[…]
@@
where $K=\mathcal{K}(I,I)$ is the kernel matrix at $I$, i.e., the kernel function $\mathcal{K}(\cdot,\cdot)$ evaluated at the discretization time points $I$; $\dot{\mathcal{X}}_{G}(I):=f(x(I),\theta,I)$ is the right hand side of the ODE evaluated at the GP discretized sample path $x(I)$; $\dot{\mathcal{G}}(I)=\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}x(I)$ is the conditional mean of the GP gradient, conditioning on the GP; and the norm matrix $C=\mathcal{K^{\prime\prime}}(I,I)-\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}\mathcal{K^{\prime}}(I,I)$ is the conditional covariance matrix of the GP gradient conditioning on the GP. Here, $\mathcal{{}^{\prime}K}=\frac{\partial}{\partial s}\mathcal{K}(s,t)$, $\mathcal{K^{\prime}}=\frac{\partial}{\partial t}\mathcal{K}(s,t)$, and $\mathcal{K^{\prime\prime}}=\frac{\partial^{2}}{\partial s\partial t}\mathcal{K}(s,t)$; $\|v\|_{A}^{2}=v^{\intercal}Av$ denotes the quadratic form. With Equation (7), MAGI employs Monte Carlo methods to sample $\theta$ and $\bm{x}(\bm{I})$ from the joint posterior. If forecasting is needed, the discretization set $I=\{\iota_{j}\}_{j=1}^{M}$ will span into the forecasting time horizon, following the recommendation from Wong et al. (2023). Other details of MAGI, including the GP specifications, can be found in Yang et al. (2021).
To clarify, under the MAGI framework, all system parameters — including $\Theta$ and component-wise noise levels $\sigma^{2}$ — are automatically estimated. These do not require user specification. The user only needs to specify the ODEs governing the system and the noisy data as input. In practice, for most experimental settings shown in this manuscript, MAGI’s entire inference routine takes at most a few minutes to run.
2.3 A note on point estimates
With its Bayesian foundation and Monte Carlo inference framework, MAGI is inherently capable of uncertainty quantification in estimation and prediction. In contrast, PINN can only output point estimates and point predictions.

### v2

@@ Abstract
In the era of AI, neural networks have become increasingly popular for modeling, inference, and prediction, largely due to their potential for universal approximation. With the proliferation of such deep learning models, a question arises: are leaner statistical methods still relevant? To shed insight on this question, we employ the mechanistic nonlinear ordinary differential equation (ODE) inverse problem as a testbed, using the physics-informed neural network (PINN) as a representative of the deep learning paradigm and manifold-constrained Gaussian process inference (MAGI) as a representative of statistically principled methods. Through case studies involving the SEIR model from epidemiology and the Lorenz model from chaotic dynamics, we demonstrate that statistical methods are far from obsolete, especially when working with sparse and noisy observations. On tasks such as parameter inference and trajectory reconstruction, statistically principled methods consistently achieve lower bias and variance, while using far fewer parameters and requiring less hyperparameter tuning. Statistical methods can also decisively outperform deep learning models on out-of-sample future prediction, where the absence of relevant data often leads overparameterized models astray. Additionally, we find that statistically principled approaches are more robust to accumulation of numerical imprecision and can represent the underlying system more faithfully to the true governing ODEs.
@@ 1 Introduction
The advancement of deep neural network models in the last fifteen years has profoundly altered the scientific landscape of estimation, prediction and decision making, from the early success of image recognition (Krizhevsky et al. 2012, He et al. 2016), to the success of self-learning of board games (Silver et al. 2017), to machine translation (Wu et al. 2016), to generative AI (Ho et al. 2020), and to the success of protein structure prediction (Jumper et al. 2021), among many other developments. In many of these successes, there are no well-established mechanistic models to describe the underlying problem — for example, we do not fully understand how human brains translate from one language to another. As such, it is conceivable that such successes are attributable to deep neural networks’ remarkable capabilities for universal function approximation. In contrast, the hand-crafted models that existed before deep neural networks (such as n-gram models (Katz 1987, Brown et al. 1992, Bengio et al. 2000)) were too restricted to offer satisfactory approximation.
@@
How well do deep neural network models work when there are well-established mechanistic models (as in physical sciences, where decades of theoretical and experimental endeavors have yielded highly accurate mechanistic models in many cases) — in particular, how do the inference and prediction results of deep neural network models compare to more statistical approaches in the presence of reliable mechanistic models — is an interesting question. The current article focuses on this comparison.
@@
In physical and natural sciences, the mechanistic models are often described by ordinary differential equations (ODEs), which are often referred to as dynamical systems. Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo 2006, Mueller and Siltanen 2012), arises in many scientific studies, where advances in experimental and data collection techniques have allowed researchers to track dynamical systems closer to real-time (Xie 2015). Although ODE models are parametric, their inference has long challenged researchers, as most nonlinear ODEs require numerical solutions, which can be expensive for many practical applications (Mueller and Siltanen 2012).
@@
In this article, we compare the neural network approach to the statistical approach on their inference of dynamical systems. Both approaches were proposed to address the numerical challenge of (nonlinear) dynamical systems. Specifically, we compare the performance of the recently developed physics-informed neural network (PINN) (Raissi et al. (2017b, 2019), Lu et al. (2021)), which represents the neural network approach to the ODE inference problem, to that of manifold-constrained Gaussian process inference (MAGI) (Yang et al. 2021), which represents a principled Bayesian statistical approach to ODE inference. In this comparison, we investigate three different aspects of statistical inference from noisy and sparse observations: (i) how well can the methods recover the model parameters (including quantities of scientific interest defined by the parameters); (ii) how well can the methods recover the system’s time-trajectories; and (iii) how well can the methods predict the future state of the system given the observations from a short time window. In the comparison, for a comprehensive understanding we also consider the case of missing system components, where some system components are entirely unobserved—such cases often arise in real experiments due to various (technical or experimental) limitations of data collection.
@@
Our study demonstrates that MAGI is generally more robust and reliable than PINN, particularly in sparse and noisy data environments that are common in laboratory sciences. In these environments, researchers often face fewer than 100 observation data points with noise levels ranging from 5%, 15% to higher. We found that for the SEIR model, when all system components are observed, the performance of PINN and MAGI is comparable. However, when missing components are present, MAGI clearly outperforms PINN. The contrast becomes even more pronounced with the chaotic Lorenz system, where PINN routinely struggles to handle the complexity of the system, while MAGI continues to consistently produce accurate inference results. Our results indicate that in the era of deep learning, statistically principled methods are still very relevant, especially when working with sparse and noisy data. Our study underscores the importance of proper probabilistic modeling in capturing the underlying dynamics for inference and prediction when one has limited data.
@@ 1.1 Statistical inference of dynamical systems
where the vector $\bm{x}(t)$ represents the system’s state/status at time $t$, and $\bm{\theta}$ is the model parameters. For example, in neuroscience, ODEs underpin models of neuron signal transmission (FitzHugh 1961, Nagumo et al. 1962). In ecology, predator-prey population dynamics uses ODEs (Lotka 1932) to capture interactions between species over time. In systems biology, ODEs map complex gene regulatory networks, enabling insight into cellular responses (Hirata et al. 2002). In epidemiology, ODE models are used to track disease progression and evaluate public health interventions (Shaman and Karspeck 2012).
Statistical inference of dynamical systems, also known as the inverse problem (Kaipio and Somersalo 2006, Mueller and Siltanen 2012), includes estimating the model parameters, inferring/recovering the time-trajectories of the system, and making predictions of the underlying system at future times from noisy observations collected at various time points. The difficulty of inferring dynamical system models lies in the computational costs associated with numerically solving the ODEs, which has long challenged the applied mathematics community (Mueller and Siltanen 2012).
Utilizing the approximating power of neural networks, PINN was recently introduced (Raissi et al. (2017b, 2019)); it uses a neural network as a function approximator to the solution of the ODE, thus bypassing numerical integration, where the loss function of the neural network incorporates the ODE dynamics. Section 2 describes PINN in detail.
Parallel to the development of PINN, there are also efforts in the statistical community for the inference of dynamical systems. One competitive method is MAGI (Yang et al. 2021), which assigns a Gaussian process prior on the system components $\bm{x}(t)$ and explicitly conditions it to ensure that the derivatives of $\bm{x}(t)$ satisfy the ODE specification. This approach enables MAGI to also bypass the need for numerical integration. Section 2 also describes MAGI in detail.
@@ 1.2 Other approaches
When the function $f$ in the ODE (Equation (1)) is nonlinear, determining $\bm{x}(t)=\mathbb{X}_{\bm{x}_{0},\bm{\theta}}(t)$ given the initial conditions $\bm{x}_{0}$ and the parameter $\bm{\theta}$ often requires a numerical integration method, such as the fifth-order Runge–Kutta (Lapidus and Seinfeld 1971) $\hat{\mathbb{X}}^{RK5}$. Given access to numerical integration, the inverse problem of estimating $\bm{\theta}$ from an experimental/observation dataset $\{y(\tau_{i}):i=1,2,\ldots,N\}$, where $y(\tau_{i})=X(\tau_{i})+\epsilon(\tau_{i})$ represents a noisy observation of the system at time $\tau_{i}$, can be cast as an optimization problem:
@@
Before the recent development of PINN and MAGI, various attempts have been made to mitigate this computational difficulty. In statistics, an innovative contribution that uses splines as a surrogate for $\bm{x}(t)$ was introduced in Ramsay et al. (2007), which formulates a generalized profiling procedure that penalizes departure from the ODE dynamics, effectively bypassing repeated numerical integration through a frequentist smoothing spline approach. In constrained optimization, surrogate models are also proposed to bypass numerical solvers for ODE inference. Hard constraints for satisfying ODEs are imposed on collocation points — for example, Gaussian process surrogates are used as convenient interpolation tools with easy computation of the hard constraints (Raissi et al. 2017a, Spitieris and Steinsland 2023, Oates et al. 2019, Chen et al. 2021). The hard constraints, however, are artificially imposed on the Gaussian process rather than a principled Bayesian derivation through conditioning (Cockayne et al. 2017, Graepel 2003, Chen et al. 2022). Separately, Barber and Wang (2014) place a GP prior on $\bm{x}(t)$ and factorize the joint model using a Dirac delta distribution enforcing the ODE; however, this factorization is a pragmatic Bayes approximation that is mathematically incompatible with the actual Bayes law (Macdonald et al. 2015).
Other statistical methods (Müller and Timmer 2002), especially the Bayesian model calibration (Kennedy and O’Hagan 2001)), have also gained popularity (Chiachio-Ruano et al. 2021, Stuart 2010, Stuart and Teckentrup 2018). Many of these methods, however, still cannot completely bypass the numerical integration steps (Marzouk et al. 2007, Lan et al. 2016, Li and Marzouk 2014, Santner et al. 2019). Those that bypass the numerical integration are mostly two-stage methods (Bär et al. 1999, Müller and Timmer 2004, Rai and Tripathi 2019), where the ODE solution and its derivatives are first reconstructed from observation data using basis function expansion, such as a polynomial basis (Franke and Schaback 1998) or splines (Bär et al. 1999). Then, the differential equation parameters are estimated using regression from the first-step approximation results. Although simple to implement, two-stage methods are not reliable when the data is sparse and noisy (Müller and Timmer 2004): they often require ad hoc follow-up approximations, such as Picard linearization or moment-matching Gaussian distribution for non-linear equations (Magnani et al. 2022, Xun et al. 2013, Zhou et al. 2022). Rather than bypassing numerical integration, Ghosh et al. (2021) retain the ODE solver but accelerate Bayesian inference itself by applying variational optimization with the reparameterization trick, achieving significant speed-ups over MCMC.
@@ 2.1 PINN: a neural network (NN) approach
Raissi et al. (2019) suggests using a neural network (“NN”) $N(t)$ as a function approximator of the system components $\bm{x}(t)$. Their approach incorporate the ODE physics information into the neural network loss function, popularizing the concept of a physics-informed neural network (PINN):
@@
where $\dot{\mathcal{X}}_{N}(t)=f(N_{\phi}(t),\theta,t)$ approximates $dx/dt$ by plugging in the NN surrogate $N_{\phi}(t)$ in place of $\bm{x}(t)$, and the gradient $\dot{N}_{\phi}$ is computed via automatic differentiation (i.e., backpropagation). The physics loss $\frac{1}{M}\sum_{j=1}^{M}\left\|\dot{\mathcal{X}}_{N}(\iota_{j})-\dot{N}_{\phi}(\iota_{j})\right\|^{2}$ is essentially a Monte Carlo integration approximation of $\int_{t}\left\|\dot{\mathcal{X}}_{N}(t)-\dot{N}_{\phi}(t)\right\|^{2}dt$ (up to a normalizing constant) using random or pseudo-random collocation/discretization time points $I=\{\iota_{j}\}_{j=1}^{M}$. If forecasting is needed, this discretization set $I=\{\iota_{j}\}_{j=1}^{M}$ could easily span into the forecasting time horizon. $\phi$ is the over-parameterized NN nuisance parameter (e.g., the weights connecting the neurons). With such a loss function as described in Equation (3), an NN training process is carried out to obtain the optimized parameters $\theta$. As with any NN-based approach, the architecture, hyper-parameters, and training strategies of PINN require careful tuning and refinement. See Wang et al. (2023b) for a discussion on popular PINN architectures and hyper-parameters, and Rathore et al. (2024) for a recommended training schedule.
@@ 2.2 MAGI: a statistical Bayesian approach
$=$ $\text{ const}+\log\pi(\theta)+\log p\Big(X(I)\Big)+\log p\Big(y(\tau)\Big|X(I)\Big)+\log p\Big(\dot{X}(I)=f(X,\theta,I)\Big|X(I)\Big)$ (6)
$=$ $\text{ const}+\log\pi(\theta)-\frac{1}{2}\Bigg\{\underbrace{\Big\|x(I)\Big\|_{K^{-1}}^{2}}_{\text{GP prior part}}+\underbrace{\sum_{i=1}^{N}\frac{1}{\sigma^{2}}\left\|y(\tau_{i})-x(\tau_{i})\right\|^{2}}_{\text{observation error part}}+\underbrace{\Big\|\dot{\mathcal{X}}_{G}(I)-\dot{\mathcal{G}}(I)\Big\|_{C^{-1}}^{2}}_{\text{mechanistic fidelity part}}\Bigg\},$ (7)
@@
where $K=\mathcal{K}(I,I)$ is the kernel matrix at $I$, i.e., the kernel function $\mathcal{K}(\cdot,\cdot)$ evaluated at the discretization time points $I$; $\dot{\mathcal{X}}_{G}(I):=f(x(I),\theta,I)$ is the right hand side of the ODE evaluated at the GP discretized sample path $x(I)$; $\dot{\mathcal{G}}(I)=\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}x(I)$ is the conditional mean of the GP gradient, conditioning on the GP; and the norm matrix $C=\mathcal{K^{\prime\prime}}(I,I)-\mathcal{{}^{\prime}K}(I,I)\mathcal{K}(I,I)^{-1}\mathcal{K^{\prime}}(I,I)$ is the conditional covariance matrix of the GP gradient conditioning on the GP. Here, $\mathcal{{}^{\prime}K}=\frac{\partial}{\partial s}\mathcal{K}(s,t)$, $\mathcal{K^{\prime}}=\frac{\partial}{\partial t}\mathcal{K}(s,t)$, and $\mathcal{K^{\prime\prime}}=\frac{\partial^{2}}{\partial s\partial t}\mathcal{K}(s,t)$; $\|v\|_{A}^{2}=v^{\intercal}Av$ denotes the quadratic form. With Equation (7), MAGI employs Monte Carlo methods to sample $\theta$ and $\bm{x}(\bm{I})$ from the joint posterior. If forecasting is needed, the discretization set $I=\{\iota_{j}\}_{j=1}^{M}$ will span into the forecasting time horizon, following the recommendation from Wong et al. (2023). Other details of MAGI, including the GP specifications, can be found in Yang et al. (2021). When forecasting in a sensitive system (e.g., Lorenz), we propose a sequential forecasting routine described in SI Section S2 that sequentially expands the discretization set a few timesteps at a time to promote MCMC convergence.
2.3 Qualitative difference and terminology
Throughout this article, we use NN-based to refer to approaches such as standard PINNs that represent the latent solution $\mathbf{X}(t)$ with a deep neural network and estimate unknowns primarily through loss-based optimization. By statistically principled methods, we refer to approaches that specify an explicit probabilistic model for the latent trajectory and observations and carry uncertainty through inference; MAGI is an example, as it begins with a Gaussian-process prior over $\mathbf{X}(t)$ and conditions on both the observations and the differential-equation operator to obtain a posterior distribution over trajectories and parameters. In PINNs, the data-fitting and the ODE residual typically enter as penalty terms in the objective. While squared-error losses can be interpreted as corresponding to Gaussian-noise assumptions, these assumptions are usually implicit rather than part of an explicit generative model. As a result, PINNs are often presented as plug-and-play tools, whereas Bayesian formulations such as MAGI make modeling assumptions and uncertainty explicit, enabling uncertainty quantification and interpretable uncertainty propagation.
In addition, while MAGI is conceptually derived from Bayesian conditioning on the ODE (i.e., enforcing the differential equation as a strict constraint), in practice this conditioning must be implemented on a finite discretization set of time points. Likewise, PINNs enforce the governing equation through a physics-residual term evaluated on a finite set of collocation points. Thus, in the implementations, both methods incorporate ODE fidelity through discretization; the key distinction is how this ODE mechanistic information is balanced against data fit. In MAGI, this balance is handled within an explicit probabilistic model, whereas in standard PINNs it is controlled primarily through user-chosen penalty weights (e.g., $\lambda$) and training choices.
2.4 A note on point estimates and uncertainty quantification
With its Bayesian foundation and Monte Carlo inference framework, MAGI naturally provides uncertainty quantification for both parameters and trajectories, including posterior credible intervals and predictive uncertainty. In contrast, standard PINNs, as commonly used in practice, primarily return point estimates and point predictions.
This distinction already highlights a practical advantage of MAGI for scientific inference, where uncertainty bands and credible intervals are often necessary for downstream interpretation and decision-making. Since standard PINNs do not natively provide calibrated posterior uncertainty, a direct comparison of uncertainty quantification would be inherently asymmetric (and would require introducing additional, nonstandard uncertainty wrappers for PINN that bring extra modeling choices and calibration procedures). We therefore,